Part B Constants Find the electric field just outside the paint layer A charged paint is spread in a very thin uniform layer over the surface of a plastic sphere of diameter 11.0 cm, giving it a charge of -11.0 HC N/C Submit Previous Answers Request Answer X Incorrect; Try Again; 4 attempts remaining Part C Find the electric field 8.00 cm outside the surface of the paint layer A2 B--2.742 106 N/C Submit
Answer #1

B)

Q = 11 uC

Electric field, E = Q / (4 pi e0 R^2)

= (11 x 10^-6) / (4 pi x 8.85 x 10^-12 x (0.11/2)^2)

= 3.27 x 10^7 N/C

C)

r = (0.11/2) + 0.08 = 0.135 m

E = Q / (4 pi e0 r^2)

= (11 x 10^-6) / (4 x pi x 8.85 x 10^-12 x (0.135)^2)

= 5.43 x 10^6 N/C

Comment in case any doubt please rate my answer ....